# Seeing Eye to AI: Human Alignment via Gaze-Based Response Rewards for Large Language Models

### v1

@@ 1 Introduction
Recent advancements in Natural Language Processing (NLP)have led to the emergence of Large Language Models (LLMs)like GPT (OpenAI, 2023), Llama (Dubey et al., 2024), Claude (Anthropic, 2024), and Gemini (Team et al., 2024), which excel across a range of tasks. These models, often consisting of billions of parameters, are trained on massive datasets and typically require extensive fine-tuning to align their outputs with human expectations  [^1] LLMs that are trained only on extensive datasets for language modeling are referred to as “pre-trained” LLMs. Subsequent approaches, such as human alignment, are categorized as “post-training”.. Several works have focused on refining the way LLMs interpret and respond to user intent (Wang et al., 2023a), which has led to the development of novel alignment techniques. A common approach to achieving human alignment involves leveraging explicit human feedback as preference information. Currently, the most widely adopted method is Reinforcement Learning from Human Feedback (RLHF)(Ouyang et al., 2022). RLHF has been implemented in many state-of-the-art LLMs (Cui et al., 2024; OpenAI, 2023; Bai et al., 2022a), and has been shown to help align models to human instructions and mitigate the generation of toxic or harmful content (Kiegeland et al., 2024). However, a persistent challenge with this approach is the difficulty of acquiring sufficient high-quality training data (Casper et al., 2023).
@@
In this work, we propose a novel approach that incorporates Eye-tracking (ET)as an additional signal to address the challenge of human alignment. ET measures oculomotor behavior i.e. the movements and fixations of the eyes, which offers insight into visual attention and information processing (Kleinke, 1986; Land and Furneaux, 1997). This allows researchers to correlate observable eye movement patterns with underlying cognitive and perceptual processes during reading and language comprehension tasks (Kleinke, 1986; Krasich et al., 2018). Moreover, ET– unlike other (explicit) forms of feedback (e.g., questionnaire data, data annotation) – does not suffer from human biases, and offers a better temporal and spatial resolution (Zhang and Hollenstein, 2024). Several studies have shown a strong correlation between human eye movements and attention patterns in transformer-based models (Wang et al., 2024a; Bensemann et al., 2022; Sood et al., 2020a). Incorporating ET data into NLP tasks has also proven valuable, as demonstrated by numerous works (Huang et al., 2023; Khurana et al., 2023; Hollenstein et al., 2019; Yang and Hollenstein, 2023; Kiegeland et al., 2024; Deng et al., 2023a; Mathias et al., 2018; McGuire and Tomuro, 2021). Recently, Kiegeland et al. (2024) proposed the integration of ET in controlled sentiment generation to create a dataset that can be used in human alignment methods.
@@
Human alignment continues to be one of the biggest challenges in the development of LLM, with RM playing an important role in addressing this issue. This paper investigates how behavioural signals, particularly ET, can be operationalised as implicit feedback to improve human alignment (the proposed approach is shown in Figure 1). Furthermore, we explore the use of ET prediction models that can generate – automatically and with little effort – ET features in response to text input, which makes our solution not only cost-effective but also highly scalable. The main contributions of our work can be summarized as follows:
@@
  • We perform for the first time an ablation study that examines several state-of-the-art LLMs, various ET prediction models, and different methods for incorporating ET features into the RM.
  • We demonstrate experimentally substantial performance improvements with the GazeReward framework, showing accuracy gains of over 20% in RM predictions across diverse human preference datasets.
@@ 2.1 Large Language Models-Human Alignment
LLMs-Human Alignment typically involves training LLMs  [^2] Before the process of human alignment, these models are referred to as “pre-trained” LLMs. on datasets curated by humans (learning from human feedback data) (Ouyang et al., 2022). This can be achieved through Supervised Fine-Tuning (SFT), where the model is trained on pairs of prompts ($x$) and corresponding human-generated responses ($y$) (Liu et al., 2024). Alternatively, alignment can be pursued via preference optimization, using a human preference dataset that differentiates between a better response ($y_{w}$) and a worse one ($y_{l}$) for the same prompt ($x$): $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$.
@@
To this day, RLHF (Ouyang et al., 2022) remains the most popular technique used in state-of-the-art LLMs like GPT-4 (OpenAI, 2023), Claude (Bai et al., 2022a), Bard (Google, 2023), and Llama 2-Chat (Touvron et al., 2023). Different implementations of RLHF can vary in terms of data collection, training processes, and choice of RL algorithms. Typically, RLHF (Ouyang et al., 2022) involves three main steps: (1) collecting feedback, (2) training a RM based on that feedback, and (3) optimising the LLMs using RL techniques, such as Proximal Policy Optimization (PPO)Schulman et al. (2017). Since RLHF was first introduced, several advancements have been made, including fine-grained reward systems (Bai et al., 2022a; Wu et al., 2023a; Dong et al., 2023a; Wang et al., 2023c, 2024b), or replaced the original PPO algorithm with other RL techniques (Wu et al., 2023b).
@@
An alternative to RLHF is DPO (Rafailov et al., 2023), which employs an offline RL approach to optimize language models based on preference data, without the need for a separate reward model. While DPO can be used independently, it is often complementary to other training methods such as SFT or statistical rejection sampling, to further improve human alignment based on a RM (Zhao et al., 2023; Liu et al., 2024; Dubey et al., 2024). Statistical rejection sampling, also called best-of-N or top-k-over-N (Bai et al., 2022a; Touvron et al., 2023; Dubey et al., 2024) is another widely used technique. Moreover, certain methods perform human alignment without RL to avoid instabilities, and fine-tune the model on filtered samples by a RM, or other sources (Dong et al., 2023b; Yuan et al., 2023).
@@
A significant challenge in all human alignment techniques is data acquisition (Casper et al., 2023). This includes problems such as evaluator misalignment, supervision difficulties, and feedback quality issues (Casper et al., 2023). However, as AI systems continue to improve, LLMs are increasingly employed for tasks that were traditionally handled by humans, such as data annotation and generation. Unlike human feedback, AI-generated feedback offers better scalability, enabling faster and more cost-effective data collection. For example, RLAIF, introduced by Bai et al. (2022a), is a promising approach that trains reward models based on preferences generated by off-the-shelf LLMs. Variations of RLAIF have been explored in several studies (Lee et al., 2023; Jiao, 2023; Cui et al., 2024; Li et al., 2024; Yang et al., 2024). In the context of self-generating instructions, approaches like Self-Instruct (Wang et al., 2023b), Self-Refine (Madaan et al., 2023), and Self-Alignment (Li et al., 2023) demonstrate how models can autonomously generate datasets based on their learned human preferences.
@@ 2.1.1 Reward modeling
In the original implementation (Ouyang et al., 2022), the goal of RM training is to train a classifier that predicts the probability of human preference $p^{*}$ between two completions (Equation 1), modelled by a Bradley-Terry model (Bradley and Terry, 1952). The typical setup involves showing two completions, with preferences being measured using win-loss-tie outcomes or a Likert scale to capture the strength of preference (Bai et al., 2022b). The data is processed into prompt-chosen-rejected trios, where the chosen completion, $y_{w}$, is preferred over the rejected one, $y_{l}$, forming the basis for training (Ouyang et al., 2022).
@@
Feature Definition First Fixation Duration (FFD) Time spent on the initial fixation Go-Past Time (GPT) Cumulative fixation time before moving to the right Total Reading Time (TRT) Overall time spent fixating on a word Number of Fixations (nFix) Number of fixations on each word Proportion of participand (fixProp) Proportion of participants that fixated on the word
@@ 2.2 Eye-Tracking
Eye-tracking (ET)systems monitor oculomotor behavior, such as eye movements and fixations, offering valuable insights into visual attention, information processing, and expands our understanding of reading and language comprehension. (Zhang and Hollenstein, 2024). Specifically, ET data often include fixations – pauses in eye movement to focus on specific areas (Mathias et al., 2022); saccades – rapid movements between two points (McGuire and Tomuro, 2021); scanpaths – sequences of fixations that reveal saccades and regressions (Yang and Hollenstein, 2023); and other temporal and spatial gaze behavior features (Zhang and Hollenstein, 2024). Incorporating ET data into NLP tasks often involves the use of several features listed in Table 1.
While several publicly available datasets such as ZUCO (Hollenstein et al., 2020a), ZUCO2 (Hollenstein et al., 2018), PROVO (Luke and Christianson, 2018), ETSA-I (Mishra et al., 2016), ETSA-II (Mishra et al., 2018), GECO (Cop et al., 2017), GECO-MT (Colman et al., 2022) are widely used in ET research, obtaining real ET data for NLP tasks remains a challenge. This is primarily due to the cost and precision requirements of ET equipment, the unavailability of gaze data during inference, as well as privacy concerns (Khurana et al., 2023). To address these challenges, two main approaches have been proposed. The first involves integrating ET data into the model during training through methods like Multi-task learning (MTL), which eliminates the need for ET data during inference (Mishra et al., 2018; Klerke et al., 2016; Ren and Xiong, 2023; Yu et al., 2024; Deng et al., 2024). The second approach involves techniques that directly predict users’ gaze behaviour (Deng et al., 2024, 2023a; Zhang and Hollenstein, 2024; Wang et al., 2024a), creating synthetic ET data for any given text or stimulus (Deng et al., 2023b; Bolliger et al., 2023; Khurana et al., 2023; Li and Rudzicz, 2021; Hollenstein et al., 2021, 2022).
@@ 3 GazeReward: Reward modeling with ET Feedback
In this section, we discuss the proposed framework for augmenting the RM using implicit feedback derived from ET signals (Figure 2). Initially, we generate the ET features (subsection 3.1). For this step, we consider two state-of-the-art ET prediction models. Next, we combine the ET features with the text (subsection 3.2), producing different types of combined embeddings, and finally pass them as input into the RM to obtain the reward for the prompt and its corresponding response (subsection 3.3).
@@ 3.2 RM augmentation using eye-tracking features
To project these features to the model embedding size ($d$), we use a Multilayer Perceptron (MLP)ET feature projector $fp()$. The $fp()$ consists of two linear layers, two dropout layers, two Layer Normalization layers, and ReLU activation, designed for stable, non-linear ET feature representation and overfitting prevention. The model’s input dimension dynamically adjusts to accommodate the number of features used during training. The ET features projector can be formulated as $emb_{ETF}=fp(f_{et})\in\mathbb{R}^{w\times d}$ (Figure 2). This formula describes the projection of ETF features ($f_{et}$) through a function $fp$, resulting in an embedding matrix $emb_{ETF}$ with dimensions $w\times d$, where $w$ represents the number of time steps and $d$ the embedding dimension.
@@ 4.1 Experimental setup
Models. As RM base models we use the pretrained checkpoint of Hugging Face for Llama 3 8B  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B (Dubey et al., 2024), Llama 3 8B-instruct 444meta-llama/Meta-Llama-3-8B-Instruct(Dubey et al., 2024) and Mistral 7B  [^5] https://huggingface.co/mistralai/Mistral-7B-v0.3 (Jiang et al., 2023a).
@@
Training procedure.
In our implementetion, the ET prediction model remains frozen while the ET features projector is trained (Figure 2). For the RM, we fine-tune the open-source models previously introduced with QLoRA (Dettmers et al., 2023) a Parameter-Efficient Fine-Tuning (PEFT)method based on Low-Rank Adaptation (LoRA)(Hu et al., 2021), with other memory optimization techniques. We follow the training process for the RM as outlined in Touvron et al. (2023); Ouyang et al. (2022). We independently train each model on its respective dataset for two epochs, as detailed in Wang et al. (2024b). For hyperparameter tuning, we reserve 15% of each dataset for validation (shown in Table 2). The best-performing checkpoints are selected based on the lowest validation loss and used for performance evaluation. We perform a grid search to determine the optimal batch size and testing values of {8, 16, 32}. The AdamW optimizer (Loshchilov and Hutter, 2019) is used, with the Learning Rate (LR)is tuned over the range {1, 5, 10, 50}e-6, following the values reported in Touvron et al. (2023); Cui et al. (2024); Wang et al. (2024b). Additionally, we evaluate different LR schedulers: constant, linear, and cosine with a minimum LR. Further hyperparameter values and implementation details for both the RM and the ET projector can be found in Appendix C.
@@ 4.2 Results
The results of our experiments on the OASST and HelpSteer datasets, covering all possible combinations of ET features, models, and inclusion methods, as shown in Table 3 and Table 4 respectively. For the Mistral model, results for the GazeAdd method are unavailable due to the inability to map features between the ET prediction model’s tokenizer and the reward model’s tokenizer (details in Appendix A.1.3). In what follows, we discuss the main findings.
@@
baseline 64.6 diff(%) 58.6 diff(%) 62.0 diff(%)
GazeConcat $fcomb_{1}$ 68.5 6.1% 68.5 17.0% 63.9 3.1%
$fcomb_{2.5}$ 70.7 9.4% 70.4 20.3% 71.2 14.7%
$fcomb_{2.2}$ 68.5 6.1% 73.1 24.8% 72.6 17.1%
GazeAdd $fcomb_{1}$ 70.9 9.8% 71.2 21.5% -
$fcomb_{2.5}$ 51.9 -19.6% 54.3 -7.2% -
$fcomb_{2.2}$ 70.0 8.3% 69.0 17.8% -
@@
baseline 54.3 diff(%) 53.2 diff(%) 51.2535 diff(%)
GazeConcat $fcomb_{1}$ 62.8 15.7% 59.6 12.0% 63.3 23.6%
$fcomb_{2.5}$ 59.9 10.3% 53.3 0.2% 53.0 3.5%
$fcomb_{2.2}$ 59.3 9.2% 53.3 0.2% 57.1 11.5%
GazeAdd $fcomb_{1}$ 63.2 16.3% 64.6 21.3% -
$fcomb_{2.5}$ 59.9 10.3% 58.8 10.5% -
$fcomb_{2.2}$ 59.1 8.7% 59.6 12.1% -
Effect of Model Initialization. We evaluate the impact of model initialization on performance. Open-access LLMs typically come in two forms: a pre-trained version without human alignment and a final version that has undergone alignment with human feedback. Since we lack access to intermediate checkpoints, we experiment with both pre-trained models (Mistral 7B and Llama 3) and models that are already human-aligned (Llama 3 Instruct). Our goal is to confirm that our method is effective for RM initialized with both pre-trained and human-aligned checkpoints. When comparing accuracy improvements relative to the baseline (without ET features), all models show considerable gains from incorporating implicit feedback. Notably, the Llama 3 8B and Mistral 7B models, which had no prior alignment, demonstrate over 20% performance improvement from the incorporation of ET features, indicating that unaligned models benefit more from implicit feedback.
@@
Eye-tracking (ET)feature importance. Different ET features capture distinct aspects of reading behaviour and information processing, influencing model performance uniquely (Zhang and Hollenstein, 2024). Here, we examine how model performance varies when incorporating three different feature combinations generated by two different ET prediction models: $fcomb_{1}$ – TRT generated by the first ET prediction model; $fcomb_{2.5}$ – five features (nFix, FFD, GPT, TRT, fixProp) generated by the second ET prediction model; and $fcomb_{2.2}$ – TRT and FFD generated by the second ET prediction model. TRT and FFD are widely used in ET research (Huang et al., 2023; Huang and Hollenstein, 2023; Zhang and Hollenstein, 2024; Maharaj et al., 2023; Wang et al., 2022), and they have been shown to correlate with attention scores from pre-trained transformer models (Wang et al., 2024a; Bensemann et al., 2022; Sood et al., 2020a). Other studies have used gradient-based saliency methods to explore similar correlations (Hollenstein and Beinborn, 2021; Wu et al., 2024b). When comparing results, we observe that the RM benefits from implicit feedback regardless of the ET feature combination or ET prediction model used. Specifically, in most cases, $fcomb_{1}$ yields the best results, particularly with the GazeAdd method. For GazeConcat, $fcomb_{2.2}$ performs best for the OASST1 dataset, while $fcomb_{1}$ excels in HelpSteer2. We attribute the superior performance of $fcomb_{1}$ to how the ET prediction model generating the fixations was trained, including the data and preprocessing methods used (see Appendix A.1.2). Moreover, when comparing $fcomb_{2.2}$ and $fcomb_{2.5}$ – both generated by the same model – only in one case does integrating nFix, GPT, and fixProp improves performance. In some instances, using $fcomb_{2.5}$ results in worse performance than the baseline, confirming findings provided by prior studies, which suggest that features related to reading time, such as FFD and TRT, contribute most to performance gains.

### v2

@@ 1 Introduction
Recent advancements in Natural Language Processing (NLP)have led to the emergence of Large Language Models (LLMs)like GPT (OpenAI, 2023), Llama (Dubey et al., 2024), Claude (Anthropic, 2024), and Gemini (Team et al., 2024), which excel across a range of tasks. These models, often consisting of billions of parameters, are trained on massive datasets and typically require extensive fine-tuning to align their outputs with human expectations  [^1] LLMs that are trained only on extensive datasets for language modeling are referred to as “pre-trained” LLMs. Subsequent approaches, such as human alignment, are categorized as “post-training”.. Several works have focused on refining the way LLMs interpret and respond to user intent (Wang et al., 2023a), which has led to the development of novel alignment techniques. A common approach to achieving human alignment involves leveraging explicit human feedback as preference information. Currently, the most widely adopted method is Reinforcement Learning from Human Feedback (RLHF)(Ouyang et al., 2024). RLHF has been implemented in many state-of-the-art LLMs (Cui et al., 2024; OpenAI, 2023; Bai et al., 2022a), and has been shown to help align models to human instructions and mitigate the generation of toxic or harmful content (Kiegeland et al., 2024). However, a persistent challenge with this approach is the difficulty of acquiring sufficient high-quality training data (Casper et al., 2023).
@@
In this work, we propose a novel approach that incorporates Eye-tracking (ET)as an additional signal to address the challenge of human alignment. ET measures oculomotor behavior i.e. the movements and fixations of the eyes, which offers insight into visual attention and information processing (Kleinke, 1986; Land and Furneaux, 1997). This allows researchers to correlate observable eye movement patterns with underlying cognitive and perceptual processes during reading and language comprehension tasks (Kleinke, 1986; Krasich et al., 2018). Moreover, ET– unlike other (explicit) forms of feedback (e.g., questionnaire data, data annotation) – does not suffer from human biases, and offers a better temporal and spatial resolution (Zhang and Hollenstein, 2024). Several studies have shown a strong correlation between human eye movements and attention patterns in transformer-based models (Wang et al., 2024a; Bensemann et al., 2022; Sood et al., 2020). Incorporating ET data into NLP tasks has also proven valuable, as demonstrated by numerous works (Huang et al., 2023; Khurana et al., 2023; Hollenstein et al., 2019; Yang and Hollenstein, 2023; Kiegeland et al., 2024; Deng et al., 2023a; Mathias et al., 2018; McGuire and Tomuro, 2021). Recently, Kiegeland et al. (2024) proposed the integration of ET in controlled sentiment generation to create a dataset that can be used in human alignment methods.
@@
Human alignment remains one of the biggest challenges in the development of LLM, with RM playing an important role in addressing this issue. This paper investigates how behavioural signals, particularly ET, can be operationalised as implicit feedback to improve human alignment (proposed approach is shown in Figure 1). Furthermore, we explore the use of ET prediction models that can generate – automatically and with little effort – ET features in response to text input, which makes our solution not only cost-effective but also highly scalable.
Our contributions are the following:
@@
  • We perform for the first time an ablation study that examines several state-of-the-art LLMs, various ET prediction models, and methods for incorporating ET features into the RM.
  • We demonstrate experimentally substantial performance improvements with the GazeReward framework, showing accuracy gains of over 10% in RM predictions across diverse human preference datasets.
@@ 2.1 Large Language Models-Human Alignment
LLMs-Human Alignment typically involves training LLMs  [^2] Before the process of human alignment, these models are referred to as “pre-trained” LLMs. on datasets curated by humans (learning from human feedback data) (Ouyang et al., 2024). This can be achieved through Supervised Fine-Tuning (SFT), where the model is trained on pairs of prompts ($x$) and corresponding human-generated responses ($y$) (Liu et al., 2024). Alternatively, alignment can be pursued via preference optimization, using a human preference dataset that differentiates between a better response ($y_{w}$) and a worse one ($y_{l}$) for the same prompt ($x$): $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$.
@@
To this day, RLHF (Ouyang et al., 2024) remains the most popular technique used in state-of-the-art LLMs like GPT-4 (OpenAI, 2023), Claude (Bai et al., 2022a), Bard (Google, 2023), and Llama 2-Chat (Touvron et al., 2023). Different implementations of RLHF can vary in terms of data collection, training processes, and choice of RL algorithms. Typically, RLHF (Ouyang et al., 2024) involves three main steps: (1) collecting feedback, (2) training a RM based on that feedback, and (3) optimising the LLMs using RL techniques, such as Proximal Policy Optimization (PPO)Schulman et al. (2017). Since RLHF was first introduced, several advancements have been made, including fine-grained reward systems (Bai et al., 2022a; Wu et al., 2023a; Dong et al., 2023a; Wang et al., 2023c, 2024b), or replaced the original PPO algorithm with other RL techniques (Wu et al., 2023b).
@@
An alternative to RLHF is DPO (Rafailov et al., 2023), which employs an offline RL approach to optimize language models based on preference data, without the need for a separate RM. While DPO can be used independently, it is often complementary to other training methods like SFT or statistical rejection sampling, to further improve human alignment based on a RM (Zhao et al., 2023; Liu et al., 2024; Dubey et al., 2024). Statistical rejection sampling, also called best-of-N or top-k-over-N (Bai et al., 2022a; Touvron et al., 2023; Dubey et al., 2024) is another widely used technique. Moreover, certain methods perform human alignment without RL to avoid instabilities, and fine-tune the model on filtered samples by a RM, or other sources (Dong et al., 2023b; Yuan et al., 2023).
@@
A major challenge in human alignment techniques is data acquisition (Casper et al., 2023). This includes issues such as evaluator misalignment, supervision difficulties, and feedback quality (Casper et al., 2023). However, as AI systems continue to improve, LLMs are increasingly employed for tasks traditionally handled by humans, such as data annotation and generation. Unlike human feedback, AI-generated feedback offers better scalability, enabling faster and more cost-effective data collection. For example, RLAIF, introduced by Bai et al. (2022a), is a promising approach that trains reward models based on preferences generated by off-the-shelf LLMs. Variations of RLAIF have been explored in several studies (Lee et al., 2023; Jiao, 2023; Cui et al., 2024; Li et al., 2024; Yang et al., 2024). In the context of self-generating instructions, approaches like Self-Instruct (Wang et al., 2023b), Self-Refine (Madaan et al., 2023), and Self-Alignment (Li et al., 2023) demonstrate how models can autonomously generate datasets based on their learned human preferences.
@@ 2.1.1 Reward modeling
In the original implementation (Ouyang et al., 2024), the goal of RM training is to train a classifier that predicts the probability of human preference $p^{*}$ between two completions (Equation 1), modelled by a Bradley-Terry model (Bradley and Terry, 1952). The typical setup involves showing two completions, with preferences being measured using win-loss-tie outcomes or a Likert scale to capture the strength of preference (Bai et al., 2022b). The data is processed into prompt-chosen-rejected trios, where the chosen completion, $y_{w}$, is preferred over the rejected one, $y_{l}$, forming the basis for training (Ouyang et al., 2024).
@@
Feature Definition First Fixation Duration (FFD) Time spent on the initial fixation Go-Past Time (GPT) Cumulative fixation time before moving to the right Total Reading Time (TRT) Overall time spent fixating on a word Number of Fixations (nFix) Number of fixations on each word Proportion of participants (fixProp) Proportion of participants that fixated on the word
@@ 2.2 Eye-Tracking
Eye-tracking (ET)systems monitor oculomotor behavior, such as eye movements and fixations, offering valuable insights into visual attention, information processing, and expands our understanding of reading and language comprehension. (Zhang and Hollenstein, 2024). Specifically, ET data often include fixations – pauses in eye movement to focus on specific areas (Mathias et al., 2020); saccades – rapid movements between two points (McGuire and Tomuro, 2021); scanpaths – sequences of fixations that reveal saccades and regressions (Yang and Hollenstein, 2023); and other temporal and spatial gaze behavior features (Zhang and Hollenstein, 2024). Incorporating ET data into NLP tasks often involves the use of several features listed in Table 1.
While several publicly available datasets such as ZUCO (Hollenstein et al., 2020a), ZUCO2 (Hollenstein et al., 2018), PROVO (Luke and Christianson, 2018), ETSA-I (Mishra et al., 2016), ETSA-II (Mishra et al., 2018), GECO (Cop et al., 2017), GECO-MT (Colman et al., 2022) are widely used in ET research, obtaining real ET data for NLP tasks remains a challenge. This is primarily due to the cost and precision requirements of ET equipment, the unavailability of gaze data during inference, as well as privacy concerns (Khurana et al., 2023). To address these challenges, two main approaches have been proposed. The first involves integrating ET data into the model during training through methods like Multi-task learning (MTL), which eliminates the need for ET data during inference (Mishra et al., 2018; Klerke et al., 2016; Ren and Xiong, 2021; Yu et al., 2024; Deng et al., 2024). The second approach involves techniques that directly predict users’ gaze behaviour (Deng et al., 2024, 2023a; Zhang and Hollenstein, 2024; Wang et al., 2024a), creating synthetic ET data for any given text or stimulus (Deng et al., 2023b; Bolliger et al., 2023; Khurana et al., 2023; Li and Rudzicz, 2021; Hollenstein et al., 2021, 2022).
@@ 3 GazeReward: Reward modeling with ET Feedback
In this section, we discuss the proposed framework for augmenting the RM using implicit feedback derived from ET signals (Figure 2). Initially, we generate the ET features (subsection 3.1) considering two state-of-the-art ET prediction models. Next, we combine the ET features with the text (subsection 3.2), producing different types of combined embeddings, and finally pass them as input into the RM to obtain the reward for the prompt and its corresponding response (subsection 3.3).
@@ 3.2 RM augmentation using eye-tracking features
To project these features to the model embedding size ($d$), we use a Multilayer Perceptron (MLP)ET feature projector $fp()$. The $fp()$ consists of two linear layers, two dropout layers, two Layer Normalization layers, and ReLU activation, designed for stable, non-linear ET feature representation and overfitting prevention. The model’s input dimension dynamically adjusts to accommodate the number of features used during training. The ET features projector can be formulated as $emb_{ETF}=fp(f_{et})\in\mathbb{R}^{w\times d}$ (Figure 2). This formula describes the projection of ETF features ($f_{et}$) through a function $fp$, resulting in an embedding matrix $emb_{ETF}$ with dimensions $w\times d$, where $w$ represents the number of tokens and $d$ the embedding dimension.
@@ 4.1 Experimental setup
Models. As RM base models we use the pretrained checkpoint of Hugging Face (Appendix A.1.4) for Llama 3 8B, Llama 3 8B-instruct (Dubey et al., 2024) and Mistral 7B (Jiang et al., 2023).
@@
Training procedure. In our implementetion the ET modules remains frozen (Figure 2). For the RM, we fine-tune the open-source models previously introduced with QLoRA (Dettmers et al., 2023) a Parameter-Efficient Fine-Tuning (PEFT)method based on Low-Rank Adaptation (LoRA)(Hu et al., 2021), with other memory optimization techniques. We follow the training process for the RM as outlined in Touvron et al. (2023); Ouyang et al. (2024). We independently train each model on its respective dataset for two epochs, as detailed in Wang et al. (2024b). For hyperparameter tuning, we reserve 15% of each dataset for validation (shown in Table 2). The best-performing checkpoints are selected based on the lowest validation loss and used for performance evaluation. We perform a grid search to determine the optimal batch size and testing values of {8, 16, 32}. The AdamW optimizer (Loshchilov and Hutter, 2019) is used, with the Learning Rate (LR)is tuned over the range {1, 5, 10, 50}e-6, following the values reported in Touvron et al. (2023); Cui et al. (2024); Wang et al. (2024b). Additionally, we evaluate different LR schedulers: constant, linear, and cosine with a minimum LR. Further hyperparameter values and implementation details for both the RM and the ET projector can be found in Appendix C.
@@ 4.2 Results
The results of our experiments on the OASST and HelpSteer datasets, covering all possible combinations of ET features, models, and inclusion methods, as shown in Table 3 and Table 4 respectively. For the Mistral model, results for the GazeAdd method are unavailable due to the inability to map features between the ET prediction model’s tokenizer and the reward model’s tokenizer (details in Appendix A.1.3). In what follows, we present key findings based on three seeds, reporting the average, mean, and statistical significance.
@@
baseline 65.9 $\pm$ 0.5 diff (%) 65.5 $\pm$ 2.1 diff (%) 66.3 $\pm$ 0.1 diff (%)
[…]
GazeAdd $fcomb_{1}$ 68.9 $\pm$ 0.9 4.6 68.9 $\pm$ 1.0 5.3 -
[…]
@@
baseline 54.7 $\pm$ 0.7 diff (%) 53.3 $\pm$ 0.8 diff (%) 54.1 $\pm$ 0.3 diff (%)
[…]
$fcomb_{2.2}$ 60.6 $\pm$ 3.3 10.9 57.9 $\pm$ 2.0 8.6 56.0 $\pm$ 2.4 3.4
[…]
Effect of Model Initialization. We evaluate the impact of model initialization on performance. Open-access LLMs typically come in two forms: a pre-trained version without human alignment and a final version that has undergone alignment with human feedback. Since we lack access to intermediate checkpoints, we experiment with both pre-trained models (Mistral 7B and Llama 3) and models that are already human-aligned (Llama 3 Instruct). Our goal is to confirm that our method is effective for RM initialized with both pre-trained and human-aligned checkpoints. When comparing accuracy improvements relative to the baseline (without ET features), all models show considerable gains from incorporating implicit feedback. Notably, the Llama 3 8B and Mistral 7B models, which had no prior alignment, demonstrate performance improvement from the incorporation of ET features, indicating that unaligned models can benefit from implicit feedback.
@@
Eye-tracking (ET)feature importance. Different ET features capture distinct aspects of reading behaviour and information processing, influencing model performance uniquely (Zhang and Hollenstein, 2024). Here, we examine how model performance varies when incorporating three different feature combinations generated by two different ET prediction models: $fcomb_{1}$ – TRT generated by the first ET prediction model; $fcomb_{2.5}$ – five features (nFix, FFD, GPT, TRT, fixProp) generated by the second ET prediction model; and $fcomb_{2.2}$ – TRT and FFD generated by the second ET prediction model. TRT and FFD are widely used in ET research (Huang et al., 2023; Huang and Hollenstein, 2023; Zhang and Hollenstein, 2024; Maharaj et al., 2023; Wang et al., 2022), and they have been shown to correlate with attention scores from pre-trained transformer models (Wang et al., 2024a; Bensemann et al., 2022; Sood et al., 2020) and with gradient-based saliency(Hollenstein and Beinborn, 2021; Wu et al., 2024b). When comparing results, we observe that the RM benefits from implicit feedback regardless of the ET feature combination or ET prediction model used. Specifically, in most cases, $fcomb_{1}$ yields the best results, particularly with the GazeAdd method. For GazeConcat, $fcomb_{2.2}$ and $fcomb_{2.5}$ performs best in general. We attribute the superior performance of $fcomb_{1}$ to how the ET prediction model generating the fixations was trained, including the data and preprocessing methods used (see Appendix A.1.2). Moreover, when comparing $fcomb_{2.2}$ and $fcomb_{2.5}$ – both generated by the same model – only in one case does integrating nFix, GPT, and fixProp improves performance. In some instances, using $fcomb_{2.5}$ results in worse performance than the baseline, confirming findings provided by prior studies, which suggest that features related to reading time, such as FFD and TRT, contribute most to performance gains.